# Supersymmetry and Localization on Three-Dimensional Orbifolds

### v1

@@ 6 Conclusions
In this paper we considered three-dimensional ${\cal N}=2$ supersymmetric field theories compactified on a large class of three-dimensional orbifolds. Our results include the calculation of the partition function of gauge theories on $\mathbb{\Sigma}\times S^{1}$ in the presence of both anti-twist and twist for the R-symmetry background gauge field, that were anticipated in [1]. We computed the one-loop determinants via the eigenvalue pairing method as well as by means of the equivariant orbifold index theorem. We also computed both the canonical expressions and the effective corrections to Chern-Simons theories and Fayet-Iliopoulos terms, which were crucial for testing our partition functions by means of non-perturbative dualities. Furthermore, we derived by supersymmetric localization the refined partition function on the branched lens space, which is a generalization of the usual lens space by the inclusion of conical singularities, thus generalizing the partition function on the so-called branched three-sphere. We also discussed a method for computing the one-loop determinants for two-dimensional quantum field theories on the spindle, again with both twist and anti-twist, extending an approach proposed in [33]. The derivation of this result exploited the interpretation of the squashed three-sphere as a smooth non-trivial circle orbi-bundle over a spindle.
[…]
@@
On the field theory side, further investigation of supersymmetric dualities is interesting to gain a deeper comprehension of the quantum moduli space pertaining to SQFTs on orbifolds. For instance, in the study of gauge theories on manifolds, the phenomenon of partition functions factorizing into holomorphic blocks is well-known [58, 52, 44, 59]. Extending such results to gauge theories on orbifolds characterized by conical singularities, as those explored in this article, is an intriguing avenue of research. Another interesting outlook lies in generalizing the findings established for twisted gauge theories on non-compact manifolds [60, 61] to cover the case of non-compact orbifolds. Moreover, gauge theory partition functions on the two-sphere are known to provide the instanton corrections to the Kähler moduli space of Calabi-Yau manifolds, which directly correlate with Gromov-Witten invariants, as shown in [62]. Extending this understanding to incorporate two-dimensional gauge theories defined on spindles presents an intriguing area for exploration. Relatedly, it would be interesting to explore possible applications of our spindle index in the context of enumerative geometry, possibly generalizing the results of [63, 64].
@@
It would be important to generalize our results to higher-dimensional models, including either $\mathcal{N}=1$ or $\mathcal{N}=2$ four-dimensional gauge theories on the direct product of a spindle and a Riemann surface, on $\mathbb{\Sigma}\times\mathbb{\Sigma}$, $\mathbb{WCP}^{2}_{{\left[{n_{1},n_{2},n_{3}}\right]}}$ (about which some results have been anticipated in [27]), or more general toric four-dimensional orbifolds. This could be done by exploiting the cohomological formalism developed e.g. in [65]. Orbifold partition functions of five-dimensional supersymmetric theories would allow to make contact with the supergravity solutions describing D4 branes wrapped on various orbifolds constructed in [66, 67, 68, 69, 70, 71, 72] and should also be relevant for six-dimensional SCFTs [73] compactified on orbifolds. Finallly, it is natural to wonder whether enlarging to the realm of orbifolds the equivariant localization program pursued in [74, 75, 76] would allow to uncover interesting aspects of (K-theoretic) Donaldson-Thomas and Gromov-Witten invariants associated to toric geometries.
@@ Acknowledgements
The authors acknowledge partial support by the INFN. We thank L. Cassia, P. Ferrero, and F. Nieri for useful conversations. We also thank A. Zaffaroni for comments on a draft of this paper.
@@ B.1 Zeta-functions and gamma-functions with multiple arguments
Appendix C CKSE
@@ E.2 Chiral multiplet
$\delta{\widetilde{F}}$ $={\rm i}\sqrt{2}(\zeta{\widetilde{\psi}}){\left[{z-\sigma-(r-2)H}\right]}-{\rm i}\sqrt{2}{\rm D}_{\mu}{\left({\zeta\gamma^{\mu}{\widetilde{\psi}}}\right)}{\color[rgb]{1,0,0}+}2{\rm i}{\widetilde{\phi}}(\zeta\lambda)~{}.$ (E.10)

### v2

@@ 6 Conclusions
In this paper we considered three-dimensional ${\cal N}=2$ supersymmetric gauge theories compactified on a large class of three-dimensional orbifolds. Our results include the calculation of the partition function of gauge theories on $\mathbb{\Sigma}\times S^{1}$ in the presence of both anti-twist and twist for the $R$-symmetry connection, which were anticipated in [25]. We computed such one-loop determinants via the eigenvalue-pairing method and by the equivariant orbifold index theorem. We also calculated both the canonical expressions and the effective corrections to Chern-Simons theories and Fayet-Iliopoulos terms, which were crucial for testing our partition functions via non-perturbative dualities. Furthermore, we derived by supersymmetric localization the refined partition function on the branched-squashed lens space, which is a generalization of the usual lens space thanks to the inclusion of conical singularities, thus extending the partition function on branched coverings of the three-sphere.
Moreover, we computed the one-loop determinants for two-dimensional quantum field theories on the spindle, again with both twist and anti-twist. The method employed in Section 5 to derive this result exploited the interpretation of the squashed three-sphere as a smooth non-trivial circle orbi-bundle over a spindle. This approach naturally extends to the computation of one-loop determinants of four-dimensional supersymmetric theories on orbifolds by deriving them from one-loop determinants of a parent five-dimensional theory. Similarly, it holds promise for calculating the one-loop determinants of gauge theories defined on six-dimensional orbifolds through reduction from seven-dimensional counterparts [57].
The results of this paper grant several avenues for future exploration. A significant next step consists in extracting the microscopic entropy of the accelerating black holes in four-dimensional anti-de Sitter spacetime [23, 24] from the large-$N$ limit of the spindle index of ${\cal N}=2$ theories with AdS4 duals [58]. This should also match the large-$N$ predictions recently obtained from supergravity considerations in [59]. Another interesting direction is probing the large-$N$ limit of partition functions of SCFTs defined on the branched-lens space ${\mathbb{\Lambda}}$, along with the investigation of their possible holographic counterparts as asympotically locally AdS4 supergravity solutions. To this aim, a possibility is to extend the supergravity solutions constructed in [14, 15, 16, 17, 60] to encompass geometries whose conformal boundaries are non-trivial circle fibrations over a spindle. It would also be interesting to study the possible implications of our findings on the supersymmetric Renyi entropy [33] and its gravity dual [61].
@@
On the field theory side, further investigation of supersymmetric dualities is interesting to gain a deeper comprehension of the quantum moduli space pertaining to SQFTs on orbifolds. For instance, in the study of gauge theories on manifolds, the phenomenon of partition functions factorizing into holomorphic blocks is well-known [62, 53, 45, 63]. Extending such results to gauge theories on orbifolds characterized by conical singularities, as those explored in this article, is an intriguing line of research. Another interesting outlook lies in generalizing the findings established for twisted gauge theories on non-compact manifolds [64, 65] to cover the case of non-compact orbifolds. Moreover, gauge theory partition functions on the two-sphere are known to provide the instanton corrections to the Kähler moduli space of Calabi-Yau manifolds, which directly correlate with Gromov-Witten invariants, as shown in [66]. Extending this understanding to incorporate two-dimensional gauge theories defined on spindles presents an intriguing area for exploration. Relatedly, it would be interesting to explore possible applications of our spindle index in the context of enumerative geometry, possibly generalizing the results of [67, 68].
@@
It would be important to generalize our results to higher-dimensional models, including either $\mathcal{N}=1$ or $\mathcal{N}=2$ four-dimensional gauge theories on the direct product of a spindle and a Riemann surface, on $\mathbb{\Sigma}\times\mathbb{\Sigma}$, $\mathbb{WCP}^{2}_{{\left[{n_{1},n_{2},n_{3}}\right]}}$ (about which some results were anticipated in [27]), or more general toric four-dimensional orbifolds. This could be done by exploiting the cohomological formalism developed e.g. in [69]. Orbifold partition functions of five-dimensional supersymmetric theories would allow to make contact with the supergravity solutions describing D4 branes wrapped on various orbifolds constructed in [70, 71, 72, 73, 74, 75, 76] and should also be relevant for six-dimensional SCFTs [77] compactified on orbifolds. Finallly, it is natural to wonder whether enlarging to the realm of orbifolds the equivariant localization program pursued in [78, 79, 80] would allow to uncover interesting aspects of (K-theoretic) Donaldson-Thomas and Gromov-Witten invariants associated to toric geometries.
@@ Acknowledgements
The authors acknowledge partial support by the INFN. We thank E. Colombo, L. Cassia, P. Ferrero and F. Nieri for useful conversations. We also thank A. Zaffaroni for comments on a draft of this paper.
@@ B.1 Zeta-functions and gamma-functions with multiple arguments
Appendix C Conformal Killing spinor equation
@@ E.2 Chiral multiplet
$\delta{\widetilde{F}}$ $={\rm i}\sqrt{2}(\zeta{\widetilde{\psi}}){\left[{z-\sigma-(r-2)H}\right]}-{\rm i}\sqrt{2}{\rm D}_{\mu}{\left({\zeta\gamma^{\mu}{\widetilde{\psi}}}\right)}+2{\rm i}{\widetilde{\phi}}(\zeta\lambda)~{}.$ (E.10)